man has $93000 invested in stocks at 8%. How much additional money should he invest in cd's at 4% so that the average return on the two investments is 5%?
Answer by mananth(16949)

percent Principal $
8.00 93000
4.00 x
5.00 93000+x

8*93000+4x=5*(93000+x)
744000+ 4x=5*93000+5x
744000+4x=465000+5x
4x-5x=465000-744000
-1x=-279000
/-1
x=$ 279000